OpenStudy (anonymous):

Nerds are you bored? Here's a question for you: Find the value of x so that the function has the given value: f(x)=6x+9;3. When you figure this out show me step by step how to do this:) I respect Nerds:)

ganeshie8 (ganeshie8):

function = f(x) = 6x + 9

ganeshie8 (ganeshie8):

you need to find value of x, which makes the function become 3 ha ?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is 3 you x if so then... f(x)=27

jimthompson5910 (jim_thompson5910):

f(x)=6x+9 3=6x+9 keep going to solve for x

OpenStudy (anonymous):

Do you think you could explain more.....

jimthompson5910 (jim_thompson5910):

you replace f(x) with 3, then solve for x This is because you want to find an x value that will give you f(x) = 3 It's going backwards from the last problem (where you plug in some given number to get another number)

jimthompson5910 (jim_thompson5910):

Do you know how to solve for x?

OpenStudy (anonymous):

Not really, I do but I always mess up

jimthompson5910 (jim_thompson5910):

you want to isolate x, so you want x all by itself

jimthompson5910 (jim_thompson5910):

to do that, you follow PEMDAS in reverse by undoing operations eg: you undo +9 by subtracting 9 from both sides in your first step 3=6x+9 3-9=6x+9-9 -6 = 6x What's next?

OpenStudy (anonymous):

you divide -6 into 6

jimthompson5910 (jim_thompson5910):

yes good

jimthompson5910 (jim_thompson5910):

giving you x = ???

OpenStudy (anonymous):

1

jimthompson5910 (jim_thompson5910):

close

OpenStudy (anonymous):

-1

jimthompson5910 (jim_thompson5910):

good, x = -1

jimthompson5910 (jim_thompson5910):

notice this when x = -1, f(x) is... f(x)=6x+9 f(-1)=6(-1)+9 f(-1)=-6+9 f(-1)=3 and that's exactly the output we were given

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

yw